4. Jerry did a submaximal YMCA cycle test. His estimated max workload was 1565kg*m*min-1. Jerry's weight is 179lbs.

d. Determine MET*min*wk-1 if they were to use your cycle ergometer prescription 30 min per day, 4 days per week.
Answer #1

The MET (Maximal Exercise Test) is the ratio of a given metabolic work rate to a standard resting metabolic rate of 1.0·kg-1 ·h -1

It is given thjat the max workload = 1565kg*m*min-1. = 3.6Kcal/min

weight = 179lbs= 81.36kg

convert this to MET = (1.6222W/weight of man)*60

=(1.622*3.6/81.36)*60

= 0.07*60= 4.6W/m2

We get MET = 4.3

we know that  1 MET reference value of 1 kcal·kg−1·h−1,

so here the person is having 4.3MET value

if they were to use your cycle ergometer prescription 30 min per day, 4 days per week.,

Then, So if he worked at a pace equivalent to 4.3 METs for 30 minutes it would be calculated as follows:

4.3 METs x 30 minutes = 129 MET minutes

Jerry participated in this activity 4 days per week:

129 MET minutes per day X 4 days per week = 516 MET minutes per week..